A small object of uniform density Rolls up a curved surface with an initial velocity v.It reaches upto maximum height of 3v^2/4g with respect to initial object.The object is 
Ring
Disc
Solid sphere
Hollowsphere 

Grade:11

1 Answers

Arun
5 years ago
From the law of conservation of energy,
mgh = 1/2mv² + 1/2 Iω²
I = mk²(where, k= radius of gyration) and, ω = v²/r²
by solving the eq.,
v²= 2gh/1+(k²/r²)
put 'h' = 3v²/4g in this eq. and by solving, you'll get,
k²/r² = 1/2
which means, given object is a DISC